OpenStudy (bruno102):

Write an equation of a line in slope intercept form that is perpendicular to the line 2x -3y = 12 and passes through the point (2, 6). Having a lot of trouble with this.

OpenStudy (anonymous):

mmmm ok

OpenStudy (anonymous):

y=-3/2x+b 6=-3+b b=9 then y=-3/2x+9

OpenStudy (anonymous):

the new equation is y=-3/2x+9

OpenStudy (bruno102):

Thanks so much!! Thanks for showing your work too :)